# Observation of non-Hermitian Dirac cones

### v1

@@
thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
Observation of non-Hermitian Dirac cones
Xinrong Xie Interdisciplinary Center for Quantum Information, State Key Laboratory of Extreme Photonics and Instrumentation, ZJU-Hangzhou Global Scientific and Technological Innovation Center, Zhejiang University, Hangzhou 310027, China International Joint Innovation Center, The Electromagnetics Academy at Zhejiang University, Zhejiang University, Haining 314400, China Key Lab. of Advanced Micro/Nano Electronic Devices & Smart Systems of Zhejiang, Jinhua Institute of Zhejiang University, Zhejiang University, Jinhua 321099, China Shaoxing Institute of Zhejiang University, Zhejiang University, Shaoxing 312000, China    Fei Ma Interdisciplinary Center for Quantum Information, State Key Laboratory of Extreme Photonics and Instrumentation, ZJU-Hangzhou Global Scientific and Technological Innovation Center, Zhejiang University, Hangzhou 310027, China International Joint Innovation Center, The Electromagnetics Academy at Zhejiang University, Zhejiang University, Haining 314400, China Key Lab. of Advanced Micro/Nano Electronic Devices & Smart Systems of Zhejiang, Jinhua Institute of Zhejiang University, Zhejiang University, Jinhua 321099, China Shaoxing Institute of Zhejiang University, Zhejiang University, Shaoxing 312000, China    W. B. Rui Department of Physics and HKU-UCAS Joint Institute for Theoretical and Computational Physics at Hong Kong, The University of Hong Kong, Pokfulam Road, Hong Kong, China HK Institute of Quantum Science & Technology, The University of Hong Kong, Pokfulam Road, Hong Kong, China    Zhaozhen Dong Interdisciplinary Center for Quantum Information, State Key Laboratory of Extreme Photonics and Instrumentation, ZJU-Hangzhou Global Scientific and Technological Innovation Center, Zhejiang University, Hangzhou 310027, China International Joint Innovation Center, The Electromagnetics Academy at Zhejiang University, Zhejiang University, Haining 314400, China Key Lab. of Advanced Micro/Nano Electronic Devices & Smart Systems of Zhejiang, Jinhua Institute of Zhejiang University, Zhejiang University, Jinhua 321099, China Shaoxing Institute of Zhejiang University, Zhejiang University, Shaoxing 312000, China    Yulin Du Interdisciplinary Center for Quantum Information, State Key Laboratory of Extreme Photonics and Instrumentation, ZJU-Hangzhou Global Scientific and Technological Innovation Center, Zhejiang University, Hangzhou 310027, China International Joint Innovation Center, The Electromagnetics Academy at Zhejiang University, Zhejiang University, Haining 314400, China Key Lab. of Advanced Micro/Nano Electronic Devices & Smart Systems of Zhejiang, Jinhua Institute of Zhejiang University, Zhejiang University, Jinhua 321099, China Shaoxing Institute of Zhejiang University, Zhejiang University, Shaoxing 312000, China    Wentao Xie Department of Physics, The Chinese University of Hong Kong, Shatin, Hong Kong SAR, China    Y. X. Zhao Department of Physics and HKU-UCAS Joint Institute for Theoretical and Computational Physics at Hong Kong, The University of Hong Kong, Pokfulam Road, Hong Kong, China HK Institute of Quantum Science & Technology, The University of Hong Kong, Pokfulam Road, Hong Kong, China    Hongsheng Chen Interdisciplinary Center for Quantum Information, State Key Laboratory of Extreme Photonics and Instrumentation, ZJU-Hangzhou Global Scientific and Technological Innovation Center, Zhejiang University, Hangzhou 310027, China International Joint Innovation Center, The Electromagnetics Academy at Zhejiang University, Zhejiang University, Haining 314400, China Key Lab. of Advanced Micro/Nano Electronic Devices & Smart Systems of Zhejiang, Jinhua Institute of Zhejiang University, Zhejiang University, Jinhua 321099, China Shaoxing Institute of Zhejiang University, Zhejiang University, Shaoxing 312000, China    Fei Gao gaofeizju@zju.edu.cn Interdisciplinary Center for Quantum Information, State Key Laboratory of Extreme Photonics and Instrumentation, ZJU-Hangzhou Global Scientific and Technological Innovation Center, Zhejiang University, Hangzhou 310027, China International Joint Innovation Center, The Electromagnetics Academy at Zhejiang University, Zhejiang University, Haining 314400, China Key Lab. of Advanced Micro/Nano Electronic Devices & Smart Systems of Zhejiang, Jinhua Institute of Zhejiang University, Zhejiang University, Jinhua 321099, China Shaoxing Institute of Zhejiang University, Zhejiang University, Shaoxing 312000, China    Haoran Xue haoranxue@cuhk.edu.hk Department of Physics, The Chinese University of Hong Kong, Shatin, Hong Kong SAR, China
@@
Relativistic quasiparticle excitations arising from band degeneracies in crystals not only offer exciting chances to test hypotheses in particle physics but also play crucial roles in the transport and topological properties of materials and metamaterials. Quasiparticles are commonly described by low-energy Hamiltonians that are Hermitian, while non-Hermiticity is usually considered detrimental to quasiparticle physics. In this work, we show that such an assumption of Hermiticity can be lifted to bring quasiparticles into non-Hermitian systems. We propose a concrete lattice model containing two non-Hermitian Dirac cones, with one hosting amplifying Dirac quasiparticles and the other hosting decaying ones. The lifetime contrast between the Dirac cones at the two valleys imposes an ultra-strong valley selection rule not seen in any Hermitian systems: only one valley can survive in the long time limit regardless of the excitation, lattice shape and other details. This property leads to an effective parity anomaly with a single Dirac cone and offers a simple way to generate vortex states in the massive case. The non-Hermitian feature of the bulk Dirac cones can also be generalized to the boundary, giving rise to valley kink states with valley-locked lifetimes. This makes the kink states effectively unidirectional and more resistant against inter-valley scattering. All these phenomena are experimentally demonstrated in a non-Hermitian electric circuit lattice.
@@
In this work, we design and experimentally realize a non-Hermitian extension of the graphene lattice with a complex conjugate pair of Dirac cones. Unlike all previously realized Dirac cones in either Hermitian or non-Hermitian systems [3, 19], the Dirac cones in our study are intrinsically non-Hermitian. That is, the Dirac quasiparticles are governed by non-Hermitian Dirac Hamiltonians with a valley-dependent background imaginary term (see Fig. 1 and Eq. (2) below). This also differs from the trivial scenario where a uniform background loss added to a Hermitian graphene lattice induces the same imaginary term for both valleys. Given these differences, we dub the excitations around the non-Hermitian Dirac cones in our system non-Hermitian Dirac quasiparticles.
Figure 1: Model for realizing non-Hermitian Dirac quasiparticles. a-c, Upper panels: Tight-binding model of conventional graphene (a), the Haldane model (b), and this work (c). The red (blue) circles represent the sublattices with on-site energy $m$ ($-m$) and the black lines indicate the nearest-neighbor couplings with strength $t$. The Orange arrows in (b) indicate the next-nearest-neighbor (NNN) couplings $t_{2}e^{i\phi}$ ($t_{2}e^{-i\phi}$) along (opposite to) the arrow. The larger (smaller) black arrows in (c) denote the NNN couplings $t_{2}+\delta$ ($t_{2}-\delta$). Lower panels: Schematics of the Dirac cone dispersion of the graphene model (a), the Haldane model (b), and the non-Hermitian model in this work (c). For graphene, the K and K valleys are related to each other by the time-reversal symmetry (TRS). For the Haldane model, the TRS breaking leads to a single Dirac cone in the K valley. The bandwidth for the single Dirac cone is determined by the gap size of the K valley, denoted by orange dashed lines. In this work, a complex conjugate pair of Dirac cones at K and K valleys are related to one another by the TRS, with colors denoting the imaginary parts of the eigenvalues. The sharp imaginary eigenvalue contrast leads to an effective single Dirac cone with a large bandwidth. d, Complex band structure of our proposed model for $m=0$, $t=1$, $t_{2}=0.1$, and $\delta=0.1$, with the colors showing the imaginary parts of the eigenvalues.
@@
where $h_{0}^{\text{tb}}=\sum_{i=1,2,3}2[t_{2}\cos{(\textbf{k}\cdot\textbf{a}_{i})}-i\delta\sin{(\textbf{k}\cdot\textbf{a}_{i})}]$, $h_{1}^{\text{tb}}=t[1+\cos{(\textbf{k}\cdot\textbf{a}_{2})}+\cos{(\textbf{k}\cdot\textbf{a}_{3})}]$, $h_{2}^{\text{tb}}=t[\sin{(\textbf{k}\cdot\textbf{a}_{2})}-\sin{(\textbf{k}\cdot\textbf{a}_{3})}]$, $h_{3}^{\text{tb}}=m$, $\sigma_{0}$ is the identity matrix, and $\sigma_{1,2,3}$ are the Pauli matrices. Here $\textbf{k}=(k_{x},k_{y})$ is the wavevector, $a_{1}=(1,0)^{\text{T}}$, $a_{2}=(-1/2,\sqrt{3}/2)^{\text{T}}$ and $a_{3}=(-1/2,-\sqrt{3}/2)^{\text{T}}$ are lattice vectors, $t$ is the nearest-neighbor coupling, $m$ is the on-site mass detuning and $t_{2}\pm\delta$ is the non-Hermitian NNN coupling. We assume that all coupling terms are real, thus preserving the TRS.
@@
where $\textbf{q}=(q_{x},q_{y})$ is the momentum relative to the Dirac points, $v=\sqrt{3}t/2$ is the group velocity, and $\tau=\pm 1$ is the valley index. Evidently, the effective Hamiltonian is composed of a conventional Dirac Hamiltonian, as in the graphene case, and a unique valley-dependent non-Hermitian term, $\tau 3\sqrt{3}i\delta\sigma_{0}$, which is the key to the formation of the non-Hermitian Dirac cones [22]. Such a valley-dependent non-Hermitian term is enforced by the TRS, which is still preserved after introducing the non-Hermitian NNN coupling (Fig. 1c). Under the TRS, the effective Hamiltonians at two valleys are related as $TH^{\text{tb}}_{\text{K}}(\textbf{q})T^{-1}=H^{\text{tb}}_{\text{K}^{\prime}}(-\textbf{q})$. Since $T=\mathcal{K}$ is the complex conjugation operation, the eigenvalues at the two valleys must have opposite imaginary parts (see Methods for detailed symmetry analysis).
It is helpful to compare this model with the Haldane model to gain more insights into its physics. In both models, the two Dirac cones have nonidentical dispersions in either the real or imaginary part. In the Haldane model, this is caused by the nonequal Dirac masses induced by the complex NNN coupling, corresponding to a real gauge field, and an on-site mass detuning. While in our case, the imaginary eigenvalue contrasting results from the nonreciprocal NNN coupling that can be interpreted as a consequence of an imaginary gauge field [23]. Thus, our model can also exhibit single Dirac cone physics in the long time limit, using non-Hermiticity instead of TRS breaking. Moreover, the bandwidth for having Dirac quasiparticles in a single valley is much larger in our case as the imaginary eigenvalue contrasting goes far beyond the low-energy limit (see Fig. 1c). At the boundary, our model can support topological modes protected by the valley Chern number, which also have valley-dependent imaginary eigenvalues and effectively show a unidirectional propagation akin to the chiral edge modes in the Haldane model (Figs. 4 and 5).
@@
We experimentally study this tight-binding model using an electric circuit lattice, which has been a popular platform for studying complex physical models [24, 25, 26, 27, 28, 29, 30]. Due to the size limit of the printed circuit board (PCB) fabrication, we segment the entire sample into pieces and interconnect them using connectors (see Methods for circuit details), as shown in Fig. 2a. The zoomed-in image in the left panel of Fig. 2b illustrates the designed circuit realizing one hexagonal cell with six lattice sites, and its schematic diagram is shown in the right panel. The nonreciprocal NNN couplings are achieved through nonreciprocal capacitances ($C1\pm C2$, denoted with orange arrows), which are realized by connecting a negative impedance converter with current inversion (INIC) in parallel with a capacitor $C_{1}$ (see Supplementary Information for more details) [27]. The nearest-neighbor couplings are realized via capacitors $C$ and the two sublattices are grounded with an LC resonator circuit containing a capacitor $C_{a/b}$ and an inductor $L$. Resistors $R_{s}$ are added from each node to the ground to avoid instabilities (see Methods for detailed analysis of circuit stability). In our designed circuit, values of $C_{a}$, $C_{b}$, and $R_{s}$ can be modified with switches composed of two-pin headers. Besides, additional components are added at the boundaries of the circuits for all the samples in this work, including suitable capacitors for obtaining the same diagonal elements of the open circuit Hamiltonian and absorbing resistors $R_{a}=30~{}\Omega$ to eliminate the reflections (see Supplementary Information for more details).
@@
where $\textbf{V}=[v_{a},v_{b}]^{\text{T}}$ are the node voltages on the two sublattices and $\omega$ is the angular frequency. $H^{\text{c}}(\textbf{k})=\sum_{i=0,1,2,3}h^{\text{c}}_{i}(\textbf{k})\sigma_{i}$ is the Hamiltonian of the circuit, with $h_{0}^{\text{c}}=\sum_{i=1,2,3}-2[C_{1}\cos{(\textbf{k}\cdot\textbf{a}_{i})}-iC_{2}\sin{(\textbf{k}\cdot\textbf{a}_{i})}]+3C+6C_{1}+(C_{a}+C_{b})/2$, $h_{1}^{\text{c}}=-C[1+\cos{(\textbf{k}\cdot\textbf{a}_{2})}+\cos{(\textbf{k}\cdot\textbf{a}_{3})}]$, $h_{2}^{\text{c}}=-C[\sin{(\textbf{k}\cdot\textbf{a}_{2})}-\sin{(\textbf{k}\cdot\textbf{a}_{3})}]$, and $h_{3}^{\text{c}}=(C_{a}-C_{b})/2$. $H^{\text{c}}$ has a similar form to the tight-binding Hamiltonian (see Eq. (1)), except that $H^{\text{c}}$ contains a global offset ($3C+6C_{1}$), which only globally shifts the eigenvalues along the real axis but has no influences on the eigenstates. Accordingly, the eigenvalues in the tight-binding model can be mapped to the frequency regime through $E(\omega)=1/\omega^{2}L-i/\omega R_{s}$.
Figure 4: Topological valley kink states with valley-dependent imaginary eigenfrequencies. a, Calculated dispersion of a heterostructure with a zigzag interface constructed by a lattice with $C_{a}=200~{}$nF and $C_{b}=100~{}$nF at the upper domain and a lattice with $C_{a}=100~{}$nF and $C_{b}=200~{}$nF at the lower domain. Other parameters are $C=100~{}$nF, $C_{1}=C_{2}=10~{}$nF, $L=1.043~{}\mu$H and $R_{s}=30~{}\Omega$. The colors denote the imaginary part of eigenfrequencies. b, Measured voltage field distributions at 222 kHz with excitations at the left (upper panel) and right (lower panel) sides. The black star denotes the source and the dashed lines mark the location of the zigzag interface. c-d, Experimentally measured edge dispersion from Fourier transform of the fields at the interface with excitations at the left (c) and right (d) sides, respectively. The colors represent the normalized Fourier intensity and the dashed lines represent the theoretical results.
@@
We then introduce an on-site mass detuning (i.e., $C_{a}=100~{}$nF and $C_{b}=200~{}$nF) to study the physics in the massive case. This results in a bandgap (Fig. 3a), and opposite orbital magnetic moment and Berry curvature at the two valleys [8, 9] (see Supplementary Information for more details), in addition to the original imaginary eigenfrequency contrasting. We excite the circuit at sublattices A and B, respectively, to measure the upper and lower bands, as given in Fig. 3b-d. It is observed that wherever the circuit is excited at sublattice A or B, only the dispersion at the K valley can be seen, similar to the massless case. Note that the upper band is also excited when we excite at sublattice B since the upper band has a larger imaginary part than the lower band (see Fig. 3a).
@@
Figure 5: Suppressed inter-valley scattering at the armchair interface. a, Plots of the real (blue dots) and imaginary (red dots) parts of the valley kink states’ eigenfrequencies at $k_{y}=0$ of a heterostructure with an armchair interface against the ratio between the nonreciprocal part $C_{2}$ and the reciprocal part $C_{1}$ of the NNN coupling. The blue and red regions denote the parity-time symmetric and parity-time symmetry-broken phases, respectively. b-c, Calculated dispersions for the two cases denoted by the black stars in (a) ($C_{2}/C_{1}=0.1$ for (b) and $C_{2}/C_{1}=1$ for (c)). The colors denote the imaginary part of the eigenfrequencies. Circuit parameters are $C=100~{}$nF, $C_{1}=10~{}$nF, $L=1.043~{}\mu$H, and $R_{s}=30~{}\Omega$. d-e, Experimentally measured edge dispersions from Fourier transform of the fields at the interface with excitations at the bottom (d) and top (e) sides, respectively. The colors represent the normalized Fourier intensity and the dashed lines represent the theoretical results. f, Measured voltage field distributions at 222 kHz with excitations at the bottom (left panel) and top (right panel). The black star denotes the source and the dashed lines mark the location of the armchair interface.
@@
When a source is attached to the left side, the kink state is excited and propagates along the interface (Fig. 4b, upper panel). However, when the source is relocated to the right end, the field is localized around the excitation without any propagation signature (Fig. 4b, lower panel). The corresponding Fourier spectra for these two cases are given in Fig. 4c and  4d, showing that the right-moving states are successfully excited while the left-moving ones are not due to their significant decay. This unidirectional propagation behavior is uniquely caused by the non-Hermiticity and is essentially different from the conventional propagation of chiral edge states induced by a nonzero Chern number. From the non-Hermitian topology point of view, this interface mimics the classic Hatano-Nelson model with a nontrivial point-gap topology [23] (see Supplementary Information for more discussions).
@@
Finally, we demonstrate that the non-Hermitian valley kink states are more robust than their Hermitian counterparts. This is basically because the difference in the imaginary eigenfrequencies of the kink states in the two valleys makes it harder for inter-valley scattering to occur. To see this, we consider the armchair interface on which the projections of the two valleys overlap. In the Hermitian case, the valley kink states are usually gapped due to the inter-valley coupling [36]. However, the imaginary eigenfrequencies contrast in our model can make the kink states decouple and thus restore gapless transport. Fig. 5a plots the evolution of the eigenfrequencies of the kink states at $k_{y}=0$ (where the kink states cross each other) as a function of the dimensionless non-Hermitian parameter $C_{2}/C_{1}$. As can be seen, a gap is present in the Hermitian case (i.e., $C_{2}=0$). As $C_{2}/C_{1}$ increases, the gap size decreases and eventually becomes zero after a parity-time phase transition. Fig. 5b and 5c show the calculated dispersions before and after the phase transition, respectively. It can be clearly seen that the edge states after the phase transition are gapless and have contrasting imaginary eigenfrequencies, similar to the zigzag case. These results are consistent with the calculations from the tight-binding model (see Supplementary Information for more details). The measured (colormaps) and calculated (dashed lines) projected dispersions for the armchair valley kink states are shown in Fig. 5d and 5e, corresponding to the excitations at the bottom and top, respectively. Again, a unidirectional propagation is observed, where the kink states can only propagate upwards but not downwards (Fig. 5f).

### v2

# Non-Hermitian Dirac cones with valley-dependent lifetimes

@@
Abstract
@@
Relativistic quasiparticles emerging from band degeneracies in crystals play crucial roles in the transport and topological properties of materials and metamaterials. Quasiparticles are commonly described by Hermitian Hamiltonians, with non-Hermiticity usually considered detrimental. In this work, we show that such an assumption of Hermiticity can be lifted to bring quasiparticles into non-Hermitian regime. We propose a concrete lattice model containing two Dirac cones with valley-dependent lifetimes. The lifetime contrast enables an ultra-strong valley selection rule: only one valley can survive in the long-time limit regardless of the excitation, lattice shape and other details. This property leads to an effective parity anomaly with a single Dirac cone and offers a simple way to generate vortex states. Additionally, extending non-Hermitian features to boundaries generates valley kink states with valley-locked lifetimes, making them effectively unidirectional and more resistant against inter-valley scattering. All these phenomena are experimentally demonstrated in a non-Hermitian electric circuit lattice.
Introduction
@@
In this work, we design and experimentally realize a non-Hermitian extension of the graphene lattice with a complex conjugate pair of Dirac cones. Unlike all previously realized Dirac cones in either Hermitian or non-Hermitian systems [3, 19], the Dirac cones in our study are intrinsically non-Hermitian. That is, the Dirac quasiparticles are governed by non-Hermitian Dirac Hamiltonians with a valley-dependent background imaginary term (see Fig. 1 and Eq. (2) below).
This also differs from the trivial scenario where a uniform background loss added to a Hermitian graphene lattice induces the same imaginary term for both valleys. Given these differences, we dub the excitations around the non-Hermitian Dirac cones in our system non-Hermitian Dirac quasiparticles.
Figure 1: Model for realizing non-Hermitian Dirac quasiparticles. a-c, Upper panels: Tight-binding model of conventional graphene (a), the Haldane model (b), and this work (c). The red (blue) circles represent the sublattices with on-site energy $m$ ($-m$) and the black lines indicate the nearest-neighbor couplings with strength $t$. The Orange arrows in (b) indicate the next-nearest-neighbor (NNN) couplings $t_{2}e^{i\phi}$ ($t_{2}e^{-i\phi}$) along (opposite to) the arrow. The larger (smaller) black arrows in (c) denote the NNN couplings $t_{2}+\delta$ ($t_{2}-\delta$). Lower panels: Schematics of the Dirac cone dispersion of the graphene model (a), the Haldane model (b), and the non-Hermitian model in this work (c). For graphene, the K and K valleys are related to each other by the time-reversal symmetry (TRS). For the Haldane model, the TRS breaking leads to a single Dirac cone in the K valley in the energy range within the bandgap of the K valley, as denoted by the orange dashed lines. In this work, a complex conjugate pair of Dirac cones at the K and K valleys are related to one another by the TRS, with colors denoting the imaginary parts of the eigenvalues. In the long time limit, only states at the K valley will survive due to the sharp contrast in lifetime between the two valleys, leading to an effective single Dirac cone behavior over a large energy range as indicated by the orange dashed lines. d, Complex band structure of our proposed model for $m=0$, $t=1$, $t_{2}=0.1$, and $\delta=0.1$, with the colors showing the imaginary parts of the eigenvalues.
@@
Results
Model for realizing non-Hermitian Dirac cones
@@ Results
where $h_{0}^{\text{tb}}=\sum_{i=1,2,3}2[t_{2}\cos{(\textbf{k}\cdot\textbf{a}_{i})}-i\delta\sin{(\textbf{k}\cdot\textbf{a}_{i})}]$, $h_{1}^{\text{tb}}=t[1+\cos{(\textbf{k}\cdot\textbf{a}_{2})}+\cos{(\textbf{k}\cdot\textbf{a}_{3})}]$, $h_{2}^{\text{tb}}=t[\sin{(\textbf{k}\cdot\textbf{a}_{2})}-\sin{(\textbf{k}\cdot\textbf{a}_{3})}]$, $h_{3}^{\text{tb}}=m$, $\sigma_{0}$ is the identity matrix, and $\sigma_{1,2,3}$ are the Pauli matrices. Here $\textbf{k}=(k_{x},k_{y})$ is the wavevector, $\textbf{a}_{1}=(1,0)^{\text{T}}$, $\textbf{a}_{2}=(-1/2,\sqrt{3}/2)^{\text{T}}$ and $\textbf{a}_{3}=(-1/2,-\sqrt{3}/2)^{\text{T}}$ are lattice vectors, $t$ is the nearest-neighbor coupling, $m$ is the on-site mass detuning and $t_{2}\pm\delta$ is the non-Hermitian NNN coupling. We assume that all coupling terms are real, thus preserving the TRS.
@@
where $\textbf{q}=(q_{x},q_{y})$ is the momentum relative to the Dirac points, $v=\sqrt{3}t/2$ is the group velocity, and $\tau=\pm 1$ is the valley index. Evidently, the effective Hamiltonian is composed of a conventional Dirac Hamiltonian, as in the graphene case, and a unique valley-dependent non-Hermitian term, $\tau 3\sqrt{3}i\delta\sigma_{0}$, which is the key to the formation of the non-Hermitian Dirac cones [22]. Note that the valley-dependent non-Hermitian term is enforced by the TRS, which is still preserved after introducing the non-Hermitian NNN coupling (Fig. 1c). Under the TRS, the effective Hamiltonians at two valleys are related as $TH^{\text{tb}}_{\text{K}}(\textbf{q})T^{-1}=H^{\text{tb}}_{\text{K}^{\prime}}(-\textbf{q})$. Since $T=\mathcal{K}$ is the complex conjugation operation, the eigenvalues at the two valleys must have opposite imaginary parts (see Methods for detailed symmetry analysis).
It is helpful to compare this model with the Haldane model to gain more insights into its physics. In both models, the two Dirac cones have nonidentical dispersions in either the real or imaginary part. In the Haldane model, this is caused by the nonequal Dirac masses induced by the complex NNN coupling, corresponding to a real gauge field, and an on-site mass detuning. While in our case, the imaginary eigenvalue contrasting results from the nonreciprocal NNN coupling that can be interpreted as a consequence of an imaginary gauge field [23]. Thus, our model can also exhibit single Dirac cone physics in the long time limit, using non-Hermiticity instead of TRS breaking. Moreover, the energy range for having states in a single valley is much larger in our case as the imaginary eigenvalue contrasting goes far beyond the low-energy limit (see Fig. 1c). At the boundary, our model can support topological modes protected by the valley Chern number, which also have valley-dependent imaginary eigenvalues and effectively show a unidirectional propagation akin to the chiral edge modes in the Haldane model (Figs. 4 and 5).
@@
Observation of massless non-Hermitian Dirac cones
We experimentally study this tight-binding model using an electric circuit lattice, which has been a popular platform for studying complex physical models because of the wide variety of electrical elements and highly flexible connectivity it offers [24, 25, 26, 27, 28, 29, 30]. Due to the size limit of the printed circuit board (PCB) fabrication, we segment the entire sample into pieces and interconnect them using connectors (see Methods for circuit details), as shown in Fig. 2a. The zoomed-in image in the left panel of Fig. 2b illustrates the designed circuit realizing one hexagonal cell with six lattice sites, and its schematic diagram is shown in the right panel. The nonreciprocal NNN couplings are achieved through nonreciprocal capacitances ($C_{1}\pm C_{2}$, denoted with orange arrows), which are realized by connecting a negative impedance converter with current inversion (INIC) in parallel with a capacitor $C_{1}$ (see Supplementary Information for more details) [27]. The nearest-neighbor couplings are realized via capacitors $C$ and the two sublattices are grounded with an LC resonator circuit containing a capacitor $C_{a/b}$ and an inductor $L$. Resistors $R_{s}$ are added from each node to the ground to avoid instabilities (see Methods for detailed analysis of circuit stability). In our designed circuit, values of $C_{a}$, $C_{b}$, and $R_{s}$ can be modified with switches composed of two-pin headers. Besides, additional components are added at the boundaries of the circuits for all the samples in this work, including suitable capacitors for obtaining the same diagonal elements of the open circuit Hamiltonian and absorbing resistors $R_{a}=30~{}\Omega$ to eliminate the reflections (see Supplementary Information for more details). With the absorbing resistors, the non-Hermitian skin effect induced by the boundaries is suppressed and we will work on the conventional Brillouin zone throughout this work.
@@
where $\textbf{V}=[v_{a},v_{b}]^{\text{T}}$ are the node voltages on the two sublattices and $\omega$ is the angular frequency. $H^{\text{c}}(\textbf{k})=\sum_{i=0,1,2,3}h^{\text{c}}_{i}(\textbf{k})\sigma_{i}$ is the Hamiltonian of the circuit, with $h_{0}^{\text{c}}=\sum_{i=1,2,3}-2[C_{1}\cos{(\textbf{k}\cdot\textbf{a}_{i})}-iC_{2}\sin{(\textbf{k}\cdot\textbf{a}_{i})}]+3C+6C_{1}+(C_{a}+C_{b})/2$, $h_{1}^{\text{c}}=-C[1+\cos{(\textbf{k}\cdot\textbf{a}_{2})}+\cos{(\textbf{k}\cdot\textbf{a}_{3})}]$, $h_{2}^{\text{c}}=-C[\sin{(\textbf{k}\cdot\textbf{a}_{2})}-\sin{(\textbf{k}\cdot\textbf{a}_{3})}]$, and $h_{3}^{\text{c}}=(C_{a}-C_{b})/2$. $H^{\text{c}}$ has a similar form to the tight-binding Hamiltonian (see Eq. (1)), except that $H^{\text{c}}$ contains a global offset ($3C+6C_{1}$), which only globally shifts the eigenvalues along the real axis but has no influences on the eigenstates. The circuit dispersion (i.e., the relationship between frequency and momentum) can be obtained by first solving Eq. (3) and then mapping the eigenvalue $E(\omega)$ to the frequency regime through $E(\omega)=1/\omega^{2}L-i/\omega R_{s}$.
Figure 4: Topological valley kink states with valley-dependent imaginary eigenfrequencies. a, Calculated dispersion of a heterostructure with a zigzag interface constructed by a lattice with $C_{a}=200~{}$nF and $C_{b}=100~{}$nF at the upper domain and a lattice with $C_{a}=100~{}$nF and $C_{b}=200~{}$nF at the lower domain. Other parameters are $C=100~{}$nF, $C_{1}=C_{2}=10~{}$nF, $L=1.043~{}\mu$H and $R_{s}=30~{}\Omega$. The colors denote the imaginary part of eigenfrequencies. b, Measured voltage field distributions at 215 kHz with excitations at the left (upper panel) and right (lower panel) sides. The black star denotes the source and the dashed lines mark the location of the zigzag interface. c-d, Experimentally measured edge dispersion from Fourier transform of the fields at the interface with excitations at the left (c) and right (d) sides, respectively. The colors represent the normalized Fourier intensity and the dashed lines represent the theoretical results.
@@
Observation of massive non-Hermitian Dirac cones
We then introduce an on-site mass detuning (i.e., $C_{a}=100~{}$nF and $C_{b}=200~{}$nF) to study the physics in the massive case. This results in a bandgap (Fig. 3a), and opposite orbital magnetic moment and Berry curvature at the two valleys [8, 9] (see Supplementary Information for more details), in addition to the original imaginary eigenfrequency contrasting. We excite the circuit at sublattices A and B, respectively, to measure the upper and lower bands, as given in Fig. 3b-d. It is observed that wherever the circuit is excited at sublattice A or B, only the dispersion at the K valley can be seen, similar to the massless case. Due to the significant contrast between imaginary parts of eigenfrequencies, modes with much larger imaginary parts (i.e., those near the band edges) dominate the measured dispersion. Note that the upper band is also excited when we excite at sublattice B since the upper band has a larger imaginary part than the lower band (see Fig. 3a).
@@
Figure 5: Suppressed inter-valley scattering at the armchair interface. a, Plots of the real (blue dots) and imaginary (red dots) parts of the valley kink states’ eigenfrequencies at $k_{y}=0$ of a heterostructure with an armchair interface against the ratio between the nonreciprocal part $C_{2}$ and the reciprocal part $C_{1}$ of the NNN coupling. The blue and red regions denote the parity-time symmetric and parity-time symmetry-broken phases, respectively. b-c, Calculated dispersions for the two cases denoted by the black stars in (a) ($C_{2}/C_{1}=0.1$ for (b) and $C_{2}/C_{1}=1$ for (c)). The colors denote the imaginary part of the eigenfrequencies. Circuit parameters are $C=100~{}$nF, $C_{1}=10~{}$nF, $L=1.043~{}\mu$H, and $R_{s}=30~{}\Omega$. d-e, Experimentally measured edge dispersions from Fourier transform of the fields at the interface with excitations at the bottom (d) and top (e) sides, respectively. The colors represent the normalized Fourier intensity and the dashed lines represent the theoretical results. f, Measured voltage field distributions at 215 kHz with excitations at the bottom (left panel) and top (right panel). The black star denotes the source and the dashed lines mark the location of the armchair interface.
Observation of non-Hermitian topological valley kink states
@@
When a source is attached to the left side, the kink state is excited and propagates along the interface (Fig. 4b, upper panel). However, when the source is relocated to the right end, the field is localized around the excitation without any propagation signature (Fig. 4b, lower panel). The corresponding Fourier spectra for these two cases are given in Fig. 4c and  4d, showing that the right-moving states are successfully excited while the left-moving ones are not due to their significant decay. This unidirectional propagation behavior is uniquely caused by the non-Hermiticity and is essentially different from the conventional propagation of chiral edge states induced by a nonzero Chern number. Besides, our non-Hermitian valley kink states are also distinct from those with on-site gain/loss that do not feature valley-dependent lifetime and thus cannot exhibit unidirectional propagation [36]. From the non-Hermitian topology point of view, this interface mimics the classic Hatano-Nelson model with a non-trivial point-gap topology [23] (see Supplementary Information for more discussions).
@@
Finally, we demonstrate that the non-Hermitian valley kink states are more robust than their Hermitian counterparts. This is basically because the difference in the imaginary eigenfrequencies of the kink states in the two valleys makes it harder for inter-valley scattering to occur. To see this, we consider the armchair interface on which the projections of the two valleys overlap. In the Hermitian case, the valley kink states are usually gapped due to the inter-valley coupling [37]. However, the imaginary eigenfrequencies contrast in our model can make the kink states decouple and thus restore gapless dispersion. Fig. 5a plots the evolution of the eigenfrequencies of the kink states at $k_{y}=0$ (where the kink states cross each other) as a function of the dimensionless non-Hermitian parameter $C_{2}/C_{1}$. As can be seen, a gap is present in the Hermitian case (i.e., $C_{2}=0$). As $C_{2}/C_{1}$ increases, the gap size decreases and eventually becomes zero after a parity-time phase transition. Fig. 5b and 5c show the calculated dispersions before and after the phase transition, respectively. It can be clearly seen that the edge states after the phase transition are gapless and have contrasting imaginary eigenfrequencies, similar to the zigzag case. These results are consistent with the calculations from the tight-binding model (see Supplementary Information for more details). The measured (colormaps) and calculated (dashed lines) projected dispersions for the armchair valley kink states are shown in Fig. 5d and 5e, corresponding to the excitations at the bottom and top, respectively. Again, a unidirectional propagation is observed, where the kink states can only propagate upwards but not downwards (Fig. 5f).
Discussion